OpenStudy (happyfeet8040):

Will give metal! The functions p(x) and g(x) are shown below: g(x) = 9x p(x) = 9x Which statement best describes the graph of p(x) and g(x)? The graphs will both have their y-intercept equal to 1. The graphs will both have their y-intercept equal to 9. The graph of g(x) will eventually exceed the graph of p(x). The graph of p(x) will eventually exceed the graph of g(x).

OpenStudy (anonymous):

Is there a typo in your equations?

OpenStudy (happyfeet8040):

No :/

OpenStudy (anonymous):

There was. You meant p(x) = 9^x. What type of equation is g(x)? What about p(x)?

OpenStudy (happyfeet8040):

Oh, my apologies.

OpenStudy (happyfeet8040):

Kinda thought it was self-explanatory.

OpenStudy (anonymous):

So what type of equation is g(x)? p(x)?

OpenStudy (happyfeet8040):

Not sure....

OpenStudy (anonymous):

g(x) is a linear equation, since as x increases, y increases proportionally to x. p(x) is an exponential equation. That means as x increases, p(x) increases exponentially. For example, if x = -1, g(-1) = -9 and p(-1) = 1/9. If x = 2, g(2) = 18 and p(2) = 81. As you can see, this means that p(x) will grow larger much faster than g(x) as x increases.

OpenStudy (anonymous):

Also, what is the value of g(0) and p(0)?

OpenStudy (happyfeet8040):

The value of p(0) is 1 I think.. Im not good in math at all sorry

OpenStudy (anonymous):

If the equation is \[p(0) = 9 ^ 0\], then what is p(0)? What is the value of any number raised to the zeroth power?

OpenStudy (happyfeet8040):

Any number raised to the power of 0 is 1

OpenStudy (anonymous):

What is g(0)? Answering this should eliminate two of your possible answers.

OpenStudy (happyfeet8040):

1?

OpenStudy (anonymous):

\[g(0) = 9 * 0\] What is g(0)?

OpenStudy (happyfeet8040):

Oh, 0?

OpenStudy (happyfeet8040):

@dcabeche

OpenStudy (anonymous):

Therefore, which two answers can we eliminate? Also, since p(x) grows much large much quicker than g(x) as x increases, which choice is the only possible answer?

OpenStudy (happyfeet8040):

C?

OpenStudy (anonymous):

D. Since p(x) grows much faster than g(x), that means that p(x) will overtake g(x).

OpenStudy (happyfeet8040):

Ohhh, okay.

OpenStudy (happyfeet8040):

I get it, thank you so much. I gave you a medal

OpenStudy (anonymous):

The equation of g(x) will look like this  And p(x) will look like this 

OpenStudy (happyfeet8040):

Ohh, okay. Thank you :)))

OpenStudy (anonymous):

No problem